# Generative Predictive Control: Flow Matching Policies for Dynamic and Difficult-to-Demonstrate Tasks

### v1

@@ III SPC is Online Generative Modeling
$\nabla_{U}\log p_{\sigma}(U\mid x)=\frac{1}{\sigma^{2}}\frac{\mathbb{E}_{\tilde{U}\sim\mathcal{N}(U,\sigma^{2})}\left[g(\tilde{U})(\tilde{U}-U)\right]}{\mathbb{E}_{\tilde{U}\sim\mathcal{N}(U,\sigma^{2})}\left[g(\tilde{U})\right]}$ (19)
@@
For simplicity of notation, we drop the conditioning on $x$ and write the target distribution as $p_{\sigma}(U)$. We also denote the normal density as $\mathcal{N}(\tilde{U};U,\sigma^{2})=q_{U}(\tilde{U})$.
The score is given by
@@
In particular, this means that the SPC update (5) is nothing more than a Monte-Carlo estimate of score ascent, e.g.,
with
$\sigma^{2}\nabla_{U}\log p_{\sigma}(U\mid x)\approx\\ \frac{\sum_{i=1}^{N}g(J(U^{(i)};x)(U^{(i)}-U)}{\sum_{i=1}^{N}g(J(U^{(i)};x)},$ (27)
where $U^{(i)}\sim\mathcal{N}(U,\sigma^{2}I)$.
The additional $\sigma^{2}$ term may seem like a troublesome annoyance, but in fact Langevin step sizes $\epsilon\propto\sigma^{2}$ are a standard recommendation in the diffusion literature [6, Algorithm 1]. Here this choice of step size emerges naturally from the SPC update (5).
This connection also sheds some light on the benefits of predictive sampling (where we merely choose the best sample). In particular, the un-noised target distribution for predictive sampling is a dirac delta concentrating all probability mass at the globally optimal solution [28, Appendix A], leading to a noised target distribution $p_{\sigma}(U\mid x)$ with a single mode at the globally optimal solution. In this sense, predicitve sampling is monte-carlo score ascent on a probability distribution with a single mode at the global optimum.
For this reason, we focus our experimental results primarily on predictive sampling. A more thorough theoretical exploration of the advantages and disadvantages of other SPC algorithms is an important topic for future work.
IV Generative Predictive Control
The previous section shows that we can think of the mean of the SPC sampling distribution, $\tilde{U}_{k}$, as a sample from a state-conditioned optimal action distribution
$\bar{U}_{k}\sim p(U\mid x_{k})\propto g(J(U;x_{k})).$ (28)
This leads to a natural question: can we train a generative model to produce $\bar{U}_{k}$? In addition to imitating the SPC update process, such a generative model
$p_{\theta}(U\mid x_{k})\approx p(U\mid x_{k}),$ (29)
parameterized by network weights $\theta$, would maintain a similar structure to the flow matching and diffusion models used in behavior cloning [1, 4, 3, 2].
Remark 2.
Observant readers might note that behavior cloning methods typically condition on observations (or a history of observations) rather than a full state estimate, which is not always readily available in practice [1]. While we use state conditioning for notational simplicity, the GPC framework also applies to observation conditioning. In fact, our implementation conditions on observations $y=h(x)$ rather than the full state $x$.
This is the basic idea behind GPC. We use data ($\bar{U}_{k},x_{k}$) from running SPC in simulation to train a flow matching model (29). This flow matching model is characterized by a time-varying state-conditional vector field
$\dot{U}=v_{\theta}(U,x,t)$ (30)
that pushes samples from $U_{t}\sim\mathcal{N}(0,I)$ at $t=0$ to the target distribution (28) at $t=1$.
The simplest approach would be to define a conditional flow matching loss analogous to (11),
$\mathcal{L}_{CFM}(\theta;U_{0},\bar{U}_{k},x_{k},t)=\\ \left\|v_{\theta}(t\bar{U}_{k}-(1-t)U_{0},x_{k},t)-(\bar{U}_{k}-U_{0})\right\|^{2},$ (31)
where $(\bar{U}_{k},x_{k})$ are data points from the SPC controller, $U_{0}\sim\mathcal{N}(0,I)$ is a sample from the proposal distribution, and $t\sim\mathcal{U}(0,1)$ is sampled uniformly along the probability flow.
Figure 2: The SPC update rule move samples $\bar{U}_{k-1}\to\bar{U}_{k}$ closer to the optimal $U^{*}$, but because the SPC update is a single score ascent step, samples do to always reach the optimal sampling distribution, especially in early iterations. To deal with this, we weigh the flow matching loss (33) according to cosine similarity with the update vector $\bar{U}_{k}-\bar{U}_{k-1}$. Shaded areas indicate regions of higher weight.
The efficiency of this approach is limited, however, by the fact that while $\bar{U}_{k}$ move in the direction of higher probability mass, $\bar{U}_{k}$ are not necessarily optimal samples from the target distribution, especially in early iterations. This is illustrated in Fig. 2 for a toy 2D example. The SPC update is a monte-carlo score ascent step, but does not necessarily reach a true sample $U^{*}$ in a single step. This means that while the flow target (arrows) are accurate in some regions (e.g., lower right corner), they may be inaccurate in other areas (e.g., upper right corner between $\bar{U}_{k}$ and $U^{*}$).
To deal with this, we propose a small modification of the flow matching loss. We leverage the cosine similarity score
@@ IV Generative Predictive Control
which measures directional similarity between vectors $\mathbf{x}$ and $\mathbf{y}$. $S_{C}$ outputs a scalar in $[-1,1]$, with high values indicating a high degree of similarity. In our case, if the update $\bar{U}_{k}-\bar{U}_{k-1}$ and the flow target $\bar{U}_{k}-U_{0}$ are close, then the sample $U_{0}\sim\mathcal{N}(0,I)$ is more informative, so we weigh it higher. In particular we define weights
$w(\bar{U}_{k},\bar{U}_{k-1},U_{0})=\\ \exp\left(-\gamma(1-S_{C}(\bar{U}_{k}-\bar{U}_{k-1},\bar{U}_{k}-U_{0})))\right),$ (33)
@@
where the hyperparameter $\gamma>0$ defines a decay rate as the two vectors move further apart. This gives us $w=1$ if $\bar{U}_{k-1}$ and $U_{0}$ are in alignment, and decays outside that range. This is illustrated with green shading in Fig. 2. The final GPC loss is given by
$\mathcal{L}_{GPC}(\theta;U_{0},\bar{U}_{k},\bar{U}_{k-1},x_{k},t)=\\ w(\bar{U}_{k},\bar{U}_{k-1},U_{0})\mathcal{L}_{CFM}(\theta;U_{0};\bar{U}_{k},x_{k},t),$ (34)
which adds a dependence on $\bar{U}_{k-1}$.
To further improve performance, GPC performs a cycle of SPC simulation and model fitting. This is illustrated in Fig. 1 and outlined in Algorithm 1.
Output: Trained flow model parameters $\theta$.
1
2while not converged do
       // Gather simulation data
3
4      for $j=1,...,N_{E}$ do
5             $x^{(j)}_{0}\sim\mathcal{X}_{0}$
6            $\bar{U}^{(j)}_{0}\sim\mathcal{N}(0,\sigma^{2}I)$
7            for $k\in[1,K]$ do
                   // Sample action sequences
8
9                  $U^{(i,j)}\sim\mathcal{N}(\bar{U}^{(j)}_{k-1},\sigma^{2}I),\quad i\in[1,N_{S}]$
                  // Parallel rollouts
11
12                  $J^{(i,j)}\leftarrow J\left(U^{(i,j)};x^{i,j}_{k-1}\right)$
                  // Update actions via SPC
13                   $\bar{U}_{k}^{(j)}\leftarrow\bar{U}_{k-1}^{(j)}+\frac{\sum_{i=1}^{N}g(J^{(i,j)})(U^{(i,j)}-\bar{U}^{(j)}_{k-1})}{{\sum_{i=1}^{N}g(J^{(i,j)})}}$
                  // Advance the simulations
14                   $x_{k}^{(j)}\leftarrow f\left(x_{k-1}^{(j)},u_{k}^{(j)}\right)$
15             end for
16
17       end for
18
      // Fit flow matching model
19       $\theta\leftarrow\arg\min_{\theta}\mathbb{E}_{t,U_{0},j,k}\left[\mathcal{L}_{GPC}\left(\theta;U_{0},\bar{U}_{k}^{(j)},\bar{U}_{k-1}^{(j)},x_{k}^{(j)},t\right)\right]$
20 end while
Algorithm 1 Generative Predictive Control
In the simulation phase, we first sample initial states $x_{0}^{(j)}$ from some set of initial conditions $\mathcal{X}_{0}$ for $N_{E}$ separate simulation environments. We then proceed to perform SPC in each environment, with $N_{S}$ of the samples coming from the typical normal proposal distribution and the remaining $N-N_{S}$ samples from the flow matching model. The SPC update uses all $N$ samples to update the action sequence $\bar{U}_{k}^{j}$ for each initial condition $j$.
When advancing the simulation (line 1), there are many valid choices for the action $u_{k}^{j}$. We could choose the initial input from $\bar{U}_{k}^{(j)}$, following the SPC scheme. Or we could choose one of the samples from $p_{\theta}$, resulting an on-policy strategy. Or we could choose some other action entirely, resulting in something akin to off-policy reinforcement learning. We found that an on-policy strategy led to more exploration of the state space in early iterations, and was generally the most effective in our examples. A more thorough evaluation of different time advancement strategies remains an important area for future research.
After collecting a dataset of states $x_{k}^{(j)}$ and action sequences $U_{k}^{(j)}$, we fit a flow matching model by minimizing $\mathcal{L}_{GPC}$ (line 1) with standard methods. The expectation in this case is over flow timesteps $t\in\mathcal{U}(0,1)$, initial samples $U_{0}\sim\mathcal{N}(0,I)$, parallel environments $j=1,\dots,N_{E}$, and simulation steps $k=1,\dots,K$.
This iterative process repeats until convergence. The inclusion of flow matching samples (line 1) enables improved performance at each iteration, while the traditional SPC samples (line 1) prevent distribution collapse.
GPC benefits from parallelism throughout Algorithm 1. In addition to parallel rollouts in the SPC update step (line 1), we can parallelize over simulation environments (line 1) and in the model training step (line 1). Our implementation [34] leverages the modern vectorization and parallelization tools in JAX [35] together with the massively parallel robotics simulation made possible by MuJoCo MJX [15].
V Using a Generative Policy
In this section we discuss two ways to use a trained GPC policy: using policy samples directly, and seeding an SPC controller.
V-A Warm-starts and Temporal Consistency
A key challenge in applying generative policies of the form $p_{\theta}(U\mid x)$ is the issue of temporal consistency [3]. In real-time feedback control, the multi-modal expressiveness of generative models presents a challenge: samples at subsequent timesteps can be drawn from different modes, leading to “jittering” behavior.
The most common solution is to merely roll out several steps of the action sequence before replanning. This forces the controller to “commit” to a particular mode, but is not suitable for highly dynamic tasks that require rapid feedback. Other alternatives include averaging over samples produced at different timesteps [3], but this does not always work well in practice [2].
$U_{0}=(1-\alpha)\mathcal{N}(0,I)+\alpha\bar{U}_{k-1}$ (35)
where $\alpha\in[0,1]$ is the warm-start level. With $\alpha=1$, the flow process is started from the previous sample $\bar{U}_{k-1}$, while $\alpha=0$ recovers the original normal proposal distribution. Because flow matching essentially defines a vector field driving samples toward a mode of the sampling distribution, flows with a high warm-start level $\alpha$ tend to stay close to the same mode as the previous sample $\bar{U}_{k-1}$.
We find that this simple warm-start procedure results in smooth and performant control signals at fast feedback rates, as shown in Sec. VII below.
V-B Seeding Sampling-based Predictive Control
Another possible use of a trained GPC policy is to add samples $\bar{U}\sim p_{\theta}(U\mid x)$ from the trained flow matching model to bootstrap an SPC controller. This also alleviates concerns with mode switching and temporal consistency, as mode switches only occur if the switch results in a lower cost, as characterized by the SPC weighting function $g(\cdot)$.
Seeding SPC with a flow matching model also alleviates many of the uni-modality concerns often raised regarding SPC algorithms [36], and can allow for effective performance with fewer samples and therefore faster control loops. This approach also opens the door to continual improvement after deployment, as these samples can be used to further refine the flow matching model as in Algorithm. 1.
We found that seeding SPC offered significantly better performance than direct flow matching on the largest examples we tested, as detailed in Sec. VII below. While SPC is computationally cheap, this improved performance does come with some limitations: a full state-estimate is required to perform the SPC rollouts, and conditioning on observations $y=h(x)$ directly is no longer feasible.
VI Domain Randomization Strategies
Domain Randomization (DR) has emerged as a key ingredient in enabling sim-to-real transfer of policies trained in simulation, particularly for reinforcement learning [37, 38]. It is reasonable to expect that DR also has a critical role to play in sim-to-real transfer of GPC policies.

### v2

@@ 3 SPC is Online Generative Modeling
$\nabla_{U}\log p_{\sigma}(U\mid x)=\frac{1}{\sigma^{2}}\frac{\mathbb{E}_{\tilde{U}\sim\mathcal{N}(U,\sigma^{2})}\left[g(\tilde{U})(\tilde{U}-U)\right]}{\mathbb{E}_{\tilde{U}\sim\mathcal{N}(U,\sigma^{2})}\left[g(\tilde{U})\right]}.$ (12)
This means that the SPC update (5) provides a Monte-Carlo estimate of score ascent, e.g.,
The additional $\sigma^{2}$ term may seem like a troublesome annoyance, but in fact Langevin step sizes $\epsilon\propto\sigma^{2}$ are a standard recommendation in the diffusion literature [11, Algorithm 1]. Here, this choice of step size emerges naturally from the SPC update (5).
This connection also sheds light on the benefits of predictive sampling (where we merely choose the best sample) [7]. In particular, the un-noised target distribution for predictive sampling is a dirac delta concentrating all probability mass at the globally optimal solution [25, Appendix A], leading to a noised target distribution $p_{\sigma}(U\mid x)$ with a single mode at the globally optimal solution. In this sense, predictive sampling approximates generative modeling for a distribution with a single mode at the global optimum. For this reason, we focus our numerical investigations primarily on predictive sampling. A more thorough theoretical exploration of the advantages and disadvantages of other SPC algorithms is an important topic for future work.
4 Generative Predictive Control
The previous section shows that we can think of the mean of the SPC sampling distribution, $\tilde{U}_{k}$, as being drawn from the state-conditioned optimal action distribution
$\bar{U}_{k}\sim p(U\mid x_{k})\propto g\left(J(U;x_{k})\right).$ (14)
This leads to a natural question: can we train a generative model to produce $\bar{U}_{k}$ directly? In addition to imitating the SPC update process, such a generative model
$p_{\theta}(U\mid x_{k})\approx p(U\mid x_{k}),$ (15)
parameterized by network weights $\theta$, would maintain a structure compatible to the flow matching and diffusion models used in behavior cloning [1, 4, 3, 2].
Remark 1.
Behavior cloning methods condition on observations $y=h(x)$ (or a history of such observations) rather than a full state estimate [1]. While we write $p_{\theta}(U\mid x)$ for notational simplicity, the GPC framework can also be applied to observation conditioning. In fact, our implementation uses observations $h(x)$ rather than the full state $x$, as detailed in Appendix E.
This is the basic idea behind GPC. We use data ($\bar{U}_{k},x_{k}$) from running SPC in simulation to train a flow matching model (15). This model is characterized by a vector field
$\dot{U}=v_{\theta}(U,x,t)$ (16)
that pushes samples from $U_{t}\sim\mathcal{N}(0,I)$ at $t=0$ to the target distribution (14) at $t=1$. To learn this vector field, we minimize a conditional flow matching loss similar to (6),
$\mathcal{L}_{GPC}(\theta;U_{0},\bar{U}_{k},x_{k},t)=\\ \left\|v_{\theta}(t\bar{U}_{k}-(1-t)U_{0},x_{k},t)-(\bar{U}_{k}-U_{0})\right\|^{2},$ (17)
where $(\bar{U}_{k},x_{k})$ are data points generated by the SPC controller, $U_{0}\sim\mathcal{N}(0,I)$ is a sample from the proposal distribution, and $t\sim\mathcal{U}(0,1)$ is sampled uniformly along the probability path.
To further improve training efficiency, we weigh data points according to cosine similarity with $\bar{U}_{k}-\bar{U}_{k-1}$. This puts greater emphasis on samples similar to $\bar{U}_{k-1}$, as illustrated by the shading in Fig. 1. Further details can be found in Appendix C.
However, directly training a generative model on SPC data is not particularly effective, as the training targets are very noisy [33]. To avoid this issue, GPC performs several cycles of SPC simulation and model fitting. This is illustrated in Fig. 1 and outlined in Algorithm 1. In each cycle, samples from the partially-trained flow matching policy are used to bootstrap SPC, providing an improved sampling distribution and thus better training data for the next model fitting step.
Output: Trained flow model parameters $\theta$.
1
2while not converged do
3       for $j=1,...,N_{E}$ do
             $x^{(j)}_{0}\sim\mathcal{X}_{0}$
              // Sample initial conditions in parallel environments
4
5            $\bar{U}^{(j)}_{0}\sim\mathcal{N}(0,\sigma^{2}I)$
6            for $k\in[1,K]$ do
7
                  $U^{(i,j)}\sim\mathcal{N}(\bar{U}^{(j)}_{k-1},\sigma^{2}I),\quad i\in[1,N_{S}]$
                    // Sample action sequences
8
                  $J^{(i,j)}\leftarrow J\left(U^{(i,j)};x^{i,j}_{k-1}\right)$
                    // Parallel rollouts
10
                  $\bar{U}_{k}^{(j)}\leftarrow\bar{U}_{k-1}^{(j)}+\frac{\sum_{i=1}^{N}g(J^{(i,j)})(U^{(i,j)}-\bar{U}^{(j)}_{k-1})}{{\sum_{i=1}^{N}g(J^{(i,j)})}}$
                    // Update actions via SPC
11
                  $x_{k}^{(j)}\leftarrow f\left(x_{k-1}^{(j)},u_{k}^{(j)}\right)$
                    // Advance the simulations
12
13             end for
14
15       end for
16
      $\theta\leftarrow\arg\min_{\theta}\mathbb{E}_{t,U_{0},j,k}\left[\mathcal{L}_{GPC}\left(\theta;U_{0},\bar{U}_{k}^{(j)},\bar{U}_{k-1}^{(j)},x_{k}^{(j)},t\right)\right]$
        // Fit flow model
17
18 end while
Algorithm 1 Generative Predictive Control
We first sample initial states $x_{0}^{(j)}$ from initial conditions $\mathcal{X}_{0}$ for $N_{E}$ parallel simulation environments. We then proceed to perform SPC in each environment, with $N_{S}$ of the samples coming from the typical Gaussian proposal distribution (line 1) and the remaining samples from the flow matching policy (line 1). The policy samples help improve performance, while the Gaussian samples prevent distribution collapse. After collecting a dataset of states $x_{k}^{(j)}$ and action sequences $U_{k}^{(j)}$, we fit the flow matching model (line 1). The expectation in this case is over flow timesteps $t\in\mathcal{U}(0,1)$, initial samples $U_{0}\sim\mathcal{N}(0,I)$, parallel environments $j=1,\dots,N_{E}$, and simulation steps $k=1,\dots,K$.
GPC benefits from parallelism throughout Algorithm 1. In addition to parallel rollouts in the SPC update step (line 1), we parallelize over simulation environments (line 1) and in the model training step (line 1). Our implementation leverages the vectorization and parallelization tools in JAX [34] together with the massively parallel robotics simulation made possible by MuJoCo MJX [8].
5 Using a Trained GPC Policy
Warm-starts: In fast feedback loops, the multi-modal expressiveness of generative models presents a challenge: samples at subsequent timesteps can be drawn from different modes, leading to “jittering” behavior. This is known as temporal consistency [3]. A common solution is to roll out several steps of the action sequence before replanning [1]. This forces the controller to “commit” to a particular mode, but is not suitable for highly dynamic tasks. Other alternatives include averaging over samples produced at different timesteps [3], but this is not always effective in practice [2].
$U_{0}=(1-\alpha)\mathcal{N}(0,I)+\alpha\bar{U}_{k-1}$ (18)
where $\alpha\in[0,1]$ is the warm-start level. With $\alpha=1$, the flow process is started from the previous sample $\bar{U}_{k-1}$, while $\alpha=0$ recovers the Gaussian proposal distribution. Because flow matching defines a vector field that drives samples toward a mode of the sampling distribution, flows with a high warm-start level $\alpha$ tend to stay close to the same mode as the previous sample, $\bar{U}_{k-1}$. We find that this simple warm-start procedure enables smooth and performant high-frequecy control.
Deploying the policy: There are two ways we might use a GPC policy. The first is to deploy the policy directly, applying actions in receeding-horizon fashion, possibly with warm-starts. We refer to this strategy as simply GPC. The second is to use policy samples for SPC, alongside samples from the usual Gaussian proposal distribution. We refer to this bootstrapping approach as GPC+. GPC+ leverages inference-time compute for better performance, but requires a state estimate from which to perform the rollouts. Ordinary GPC does not require a state estimate, as the policy can be conditioned on arbitrary observations (Remark 1).
Risk-Aware Domain Randomization: Domain (DR) has emerged as a key ingredient in enabling sim-to-real transfer of policies trained in simulation, particularly for reinforcement learning [35, 36]. It is reasonable to expect that DR will also play a critical role in sim-to-real transfer of GPC policies.
Interestingly, the advent of massively parallel simiulators and the structure of the SPC/GPC paradigm enables a range of new DR possibilities. In particular, we can consider a whole family of DR strategies ranging from simple average-case robustness to more sophisticated risk metrics like conditional value-at-risk (CVar) [37, 38]. These strategies are detailed in Appendix D.
6 Simulation Studies
Figure 2: Systems used to evaluate GPC performance in simulation, from left to right: inverted pendulum, cart-pole, double cart-pole, push-T, planar walker, luffing crane, humanoid standup.
In this section, we consider the seven test systems shown in Fig. 2. We aim to answer the following:
  1. 1.
    Can GPC perform tasks that require multi-modal reasoning, as well highly dynamic tasks that require high-frequency feedback (Sec. 6.1)?
  2. 2.
    Does GPC continually improve policy performance over multiple iterations (Sec. 6.2)?
  3. 3.
    How do different domain randomization strategies impact performance (Sec. 6.3)?
  4. 4.
    What are the scalability limits of this approach (Sec. 6.4)?
In short, GPC is effective for control systems with fast dynamics at high feedback rates, enjoys the training stability characteristic of supervised learning methods, and enables risk-aware control, but struggles to scale to our largest and most difficult example (humanoid standup). We provide further discussion of these scalability limits, and how they might be overcome, in the limitations section.
We test GPC on seven simulated systems of varying state dimension and task difficulty, shown in Fig. 2. The pendulum, cart-pole, and double cart-pole are tasked with balancing upright. In push-T, an actuated finger pushes a block to a goal pose. The walker aims to move forward at a constant velocity, while the crane swings its payload to a target position. The humanoid attempts to stand up from arbitrary intitial configurations. Further details can be found in Appendix E.
The smaller examples use a simple multi-layer perceptron for the flow network $v_{\theta}$, while the larger examples use a convolutional network with FiLM conditioning [39], as recommended in [1]. All experiments were performed on a desktop computer with an NVIDIA RTX 4070 (12 GB) GPU. When evaluating trained policies, we use MuJoCo CPU rather than MJX.
6.1 Policy Performance
Footage of closed-loop GPC performance on each of the examples is shown in the accompanying video [40]. To evaluate performance quantitatively, we perform 100 simulations of SPC, GPC, and GPC+ from random initial conditions and record the average cost per time step (Fig. 4). GPC inference times range between 1 and 10 milliseconds, enabling feedback rates between 100-1000 Hz. Fig. 4 provides some insight into relative performance, but does not capture the full picture. For instance, on the walker we find that GPC enables smoother actions with far less “stumbling” than SPC, though cost per step indicates similar performance.
We find that GPC can handle multi-modal action distributions. This is evidenced by effectiveness on the push-T, which requires multi-modal reasoning to reach around the block [1]. Interestingly, GPC training takes under 20 minutes, while training a similar diffusion policy (with full state observations) takes around an hour, not counting the time to gather demonstrations [1].
Figure 3: Closed-loop double cart-pole performance with and without warm-starts. The warm-started policy (right) produces smooth actions and is able to successfully balance. Without warm-starts (left), actions jitter between modes, and the robot fails to balance.
More importantly, GPC can control systems with fast dynamics at high control rates. The double cart-pole provides a particularly clear illustration of this fact, as well as the importance of warm-starts. Fig. 3 shows performance with and without warm-starts. Without warm-starts (left, $\alpha=0$), the control actions are dominated by significant noise (top plot) and the system is unable to swing upright (bottom plot). With warm-starts (right, $\alpha=1$), we obtain smoother control actions and the robot successfully balances around the upright configuration. The controller is still able to respond rapidly to the chaotic system dynamics, as evidenced by the rapid changes between 1 and 2 seconds.
Figure 4: Average cost per time step for SPC, GPC (policy alone), and GPC+ (policy + sampling). Black bars indicate standard deviation over 100 ten-second simulations from randomized initial conditions. Applying the GPC policy directly provides performance on-par-with or better-than SPC in all cases except humanoid standup. GPC+ meets or exceeds the performance of the other methods across all examples.
6.2 Training Stability
During training process, we find that average cost of policy samples decreases monotonically between iterations, modulo noise from initial conditions. This indicates that GPC’s cycle of training and sampling does continually improve performance. These and other training curves are available in Appendix F. While we leave a systematic hyperparameter sensitivity study for future work, our empirical observation is that GPC benefits from the training stability of supervised learning. This is in contrast to reinforcement learning methods, which can exhibit high sensitivity to reward tuning, implementation details, and even the random seed used for training [15, 14].
6.3 Risk-Aware Domain Randomization
No DR Avg. CVaR
No model error 106 103 133
With model error 165 184 139
Table 1: Time (seconds) for the crane to visit a sequence of 50 randomly generated payload targets, under policies trained without DR (“No DR”), with standard DR (“Avg.”), and with a risk-averse strategy (“CVaR”). CVaR improves robustness to model error, at the cost of worse performance under nominal conditions.
We use the luffing crane example to explore the impact of different DR strategies. Specifically, we train three GPC policies: one with no DR, one with standard average-cost DR, and one with a more conservative CVaR(0.25) strategy. We use 8 randomized domains, each with slightly different joint damping, payload mass, payload inertia, and actuator gains.
After training, we apply GPC directly with warm-starts. To evaluate closed-loop performance, we generate 50 random target locations that the payload must visit. The target is moved to the next location once the payload is moved within 15 cm or after 10 seconds, whichever comes first. Table 1 reports the total time to visit all 50 targets: lower times indicate better performance. The same sequence of target locations are used for each policy.
Without model error in the simulator, both the non-randomized policy and average-case DR perform significantly better than the more conservative CVaR policy. When we add model error to the simulator (lower joint damping, heavier payload mass), all methods perform worse. But the more conservative CVaR policy degrades the least, and significantly outperforms the others.
6.4 Scalability
In choosing our set of example systems, which range from 1 to 29 degrees-of-freedom, we aim to find the scalability limits of Algorithm 1. This scalability hits a bottleneck with the largest humanoid example: the GPC policy alone is unable to reliably stand up, though GPC+ is effective.
It is likely that further cost and hyperparameter tuning, curriculum training, and more computation could improve performance for the humanoid example. However, state-of-the-art reinforcement learning methods are able to train similar behaviors [41, 42], suggesting that further algorithmic improvements in the GPC framework are necessary to achieve reliable scalability to such systems.
7 Conclusion
We introduced generative predictive control (GPC), a framework for learning flow matching policies on dynamic tasks that are easy to simulate but difficult to demonstrate. GPC leverages tight connections between generative modeling and sampling-based predictive control to generate training data for supervised learning without expert demonstrations. We showed how warm-started GPC policies enable real-time high-frequency control, ensuring temporal consistency via warm-starts. Future work will focus on validating the GPC framework on hardware, incorporating value function learning, and training generalist multi-task policies.
Limitations
Limited effectiveness on our largest (humanoid standup) example is the most severe limitation of our proposed method, as noted above. We believe that value function learning will be a critical component in overcoming this limitation. In addition to being a key element of state-of-the-art reinforcement learning methods, value learning would allow us to reduce the planning horizon $T$ in Problem 1. Reducing the planning horizon reduces the dimensionality of the sampling space, making planning easier while maintaining long-horizon reasoning. Methods that leverage connections between the gradient of the value function and the flow field $v_{\theta}$—which is closely related to the score $\nabla\log p(U\mid x)$ and therefore the gradient of the cost (2)—are of particular interest.
For the basic GPC framework introduced here, sample complexity is relatively poor: to generate a single training data point, we run $N$ short simulations—and even more under a risk-aware DR strategy. While advances in computing speeds may mean that the benefits of a supervised learning framework eventually outweigh the computational costs, methods of more fully using all of the data from SPC rollouts could significantly improve the GPC framework.
Other performance improvements could come from more benign algorithmic details. For instance, we represent action sequences with a simple zero-order-hold spline parameterization. Higher-order splines [7] or alternative parameterizations could be more effective. Better choices of action space, such as using task-space/end-effector coordinates rather than joint coordinates, could also be useful. Actuation limits are a critical component of many robotics tasks, but our implementation does not handle these in any particularly special way. Leveraging recent advances in constrained generative modeling [43, 44, 45] to do so is another potentially fruitful area for future work.
Future work will include evaluating GPC on hardware. Hardware experiments are a critical for evaluating any robot control framework: this is particularly the case when it comes to domain randomization strategies. Hardware experiments will also provide an important platform for exploring policies conditioned on various observations, ranging from raw sensor data to images to foundation model embeddings [46]. Integrated simulation and rendering in the recently-released MuJoCo playground [47] could provide a useful software platform for training image-conditioned policies.
Finally, the GPC framework offers a path toward including dynamic and difficult-to-demonstrate tasks in a generalist policy or large behavior model that combines data from many tasks [2, 16, 17]. Exploring this possibility remains an important topic for future work.
Appendix A Appendix: Common SPC Algorithms
As discussed in Section 2.1, various SPC algorithms can be characterized by the weighting function $g(J)$. For instance, MPPI [6] uses a Bolzmann-like exponentially weighted average,
$g_{MPPI}(J)=\exp\left(-J/\lambda\right),$ (19)
where $\lambda>0$ is the temperature parameter. A smaller $\lambda$ give more weight to the lowest-cost samples.
In the low-temperature limit we recover predictive sampling [7],
$g_{PS}(J)=\lim_{\lambda\to 0}\exp\left(-J/\lambda\right),$ (20)
Another popular option is CEM [22, 5], which weighs the top performing samples equally,
Tsallis-MPPI [29] provides a middle ground between MPPI and CEM via a generalized exponential,
$g_{T}(J)=\max(1-(r-1)J/\lambda,0)^{\frac{1}{r-1}},$ (22)
where the original MPPI update is recovered as $r\to 1$.
Pratical SPC performance relies heavily on dimensionality reduction via spline representations, e.g., representing $U=[u_{0},u_{1},\dots,u_{T}]$ with $T^{\prime}\ll T$ knot points. While our implementation uses only the simplest zero-order-hold interpolation, the GPC framework is compatible with more sophisticated schemes such as Fourier features or higher-order splines [7].
Appendix B Appendix: Proof of Proposition 3
@@
For simplicity of notation, we drop the conditioning on $x$ and write the target distribution as $p_{\sigma}(U)$. We also denote the normal density as
$q_{U}(\tilde{U})\triangleq\mathcal{N}(\tilde{U};U,\sigma^{2}).$
The score of the target distribution is given by
@@
Appendix C Appendix: Cosine-Distance Objective Weighting
We use a small modification of the flow matching loss based on the cosine similarity score
@@ Appendix C Appendix: Cosine-Distance Objective Weighting
which measures directional similarity between vectors $\mathbf{x}$ and $\mathbf{y}$. $S_{C}$ outputs a scalar in $[-1,1]$, with high values indicating a high degree of similarity.
In our case, we are more confident that the flow-matching objective $\mathcal{L}_{GPC}$ (6) drives samples toward the true (but unknown) optimal action sequence $U^{*}$ if the update $\bar{U}_{k}-\bar{U}_{k-1}$ and the flow target $\bar{U}_{k}-U_{0}$ are close. This is illustrated in the “flow matching” portion of Fig. 1. Darker shading indicates values of $U_{0}$ (typically drawn from a standard Gaussian) that are more informative. In particular, samples in the upper left-hand corner could be pushed toward $\bar{U}_{k}$ but away from $U^{*}$, while samples in the lower right would flow in the correct direction.
To capture this mathematically, we define weights
$w(\bar{U}_{k},\bar{U}_{k-1},U_{0})=\exp\left(-\gamma(1-S_{C}(\bar{U}_{k}-\bar{U}_{k-1},\bar{U}_{k}-U_{0})))\right),$ (30)
@@
where the hyperparameter $\gamma>0$ defines a decay rate as the two vectors move further apart (our implementation uses $\gamma=2.0$). This gives us $w=1$ if $\bar{U}_{k-1}$ and $U_{0}$ are close in the cosine similarity sense, and $w\to 0$ as they move further apart. The final training loss is given by
$w(\bar{U}_{k},\bar{U}_{k-1},U_{0})\mathcal{L}_{GPC}(\theta;U_{0};\bar{U}_{k},x_{k},t).$ (31)
This modification is motivated by the fact that while the SPC update is a monte-carlo score ascent step, it does not obtain a true sample $U^{*}$ from the optimal target distribution. In other words, $\bar{U}_{k}$ is closer to an optimal sample than $\bar{U}_{k-1}$, but it is not the case that $\bar{U}_{k}=U^{*}$. This is an important distinction between the GPC setting and the typical generative modeling setting, where samples from the target distribution are freely available.
Appendix D Appendix: Domain Randomization Strategies